Find the indicated intercept(s) of the graph of the function. x2 + 8x 23) x-intercepts of f(x) = x2 + 5x - 5

please and thank you
Answer #1

x+88 x45x-5 y = x 788/2 + 5x-5 For x-intercepts, y=0, so we have = x²+8x=(275x-5) xo = x+8x =) x² + 5x-5 x+8x = 0 (+8)=0 x=0

answered by: ANURANJAN SARSAM